OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line? A) 4y − 5x = 5 B) y − 4x = 5 C) 5y + 4x = 5 D) 5y − 4x = 5

OpenStudy (anonymous):

(4/5 in the question)

OpenStudy (raden):

use the formula : y - y1 = m (x - x1) in the equation given (x1,y1) = (0,1) and the slope m = 4/5 plug in all value above into formula

OpenStudy (anonymous):

Thank you. How do you solve it??

OpenStudy (raden):

you just need simplify the equation or modificate it till you see the answer in options. y - y1 = m (x - x1) y - 1 = 4/5 (x - 0) y - 1 = 4/5 (x) on both sides multiply by 5, get 5(y-1) = 5 * 4/5 (x) 5y - 5 = 4x substracted by 4x on both sides, get 5y - 5 - 4x = 4x - 4x 5y - 5 - 4x = 0 adds 5 on both sides, get 5y - 5 - 4x + 5 = 0 + 5 5y - 4x = 5

OpenStudy (anonymous):

Thanks a bunch!!

OpenStudy (raden):

you're welcome ;)